OpenStudy (anonymous):

Please explain how to solve one absolute value equation by writing it as two equations.

OpenStudy (anonymous):

Well we know that an absolute value will always return an answer that is nonnegative. |x| = x if x >= 0, -x if x < 0 For example, y = |5x - 3|. If 5x - 3 >= 0 then 5x >= 3, and x >= 3/5. So whenever x >= 3/5, then y = 5x - 3. Otherwise y = -(5x -3). You can solve these by setting up two equations. One equation equal to whatever is inside the absolute value function. And another equation equal to the negative of that. If y = |5x - 3| then y = 5x - 3 or y = -5x + 3. A working example: |3x - 5| = 7 This means 3x - 5 = 7 or -(3x - 5) = 7 Solve each equation separately. 3x - 5 = 7 3x = 12 x = 4 -(3x - 5) = 7 3x - 5 = -7 3x = -2 x = -2/3 x = 4 or -2/3